OpenStudy (anonymous):

The behaviour of a wire under tensile stress may be described in terms of Young Modulus E of the material of the wire and the force per unit extension of the wire. For a wire of length L and cross-sectional area A, hat is the relation between E and k?

OpenStudy (anonymous):

a)E=A/kL b)E=kA/L c)E=kL/A d)E=L/kA

OpenStudy (anonymous):

modulus = stress/strain stress = force per unit area strain = amount of deformation -> Young's modulus is for linear deformation so, change in length

OpenStudy (anonymous):

Okay, I know that E=stress/strain. Stress=F/A and Strain= k/L. But I can't do beyond this.

OpenStudy (anonymous):

I mean how to eliminate F from the equation? :O

OpenStudy (anonymous):

he said E = force per unit extension = F/delta L

OpenStudy (anonymous):

\[ E=\frac{F}{A}\frac{L}{\Delta L}\\ F=-k\Delta L \]

OpenStudy (anonymous):

What is the negative sign for? :O

OpenStudy (anonymous):

oh, neglect that for now it means, opposition

OpenStudy (anonymous):

Umm. I did not quite get how you substituted force. Can you please explain again?

OpenStudy (anonymous):

there is only one place I know where "k" comes in... the spring.

OpenStudy (anonymous):

You mean the spring constant?

OpenStudy (anonymous):

yep